OpenStudy (anonymous):

what is the value of x in the equation 2/x -3=26/x

OpenStudy (anonymous):

2/x-3=26/x multiply out by x to give 2-3x=26 3x=24 x=8

OpenStudy (anonymous):

what do u mean multiply out by x

OpenStudy (anonymous):

\[\frac{2}{x} \times x = 2\] for example and the answer is -8 sorry I missed a negative off

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

np, did you get it :D

OpenStudy (anonymous):

yea